OpenStudy (anonymous):

The ordinate of a point P is twice the abscissa. This point is equidistant from (-3, 1) and (8,-2). Find the coordinates of P

OpenStudy (anonymous):

the point P is equal to (a,b) where b is equal to 2 * a. the point P can therefore be represented as (a,2a). the point P is equidistant from the points B(-3,1) and F(8,-2) this means that the length of the line between P and B is equal to the length of the line between P and F. the equation for the length of a line is: L = sqrt((x2-x1)^2 + (y2-y1)^2) Assign one the points for each line to (x1,y1) and assign the other point for each line to (x2,y2) for the first line, you get: (x1,y1) = (-3,1) (x2,y2) = (a,2a) for the second line, you get: (x1,y1) = (8,-2) (x2,y2) = (a,2a) the point (x2,y2) is the point that's common to both lines. for each line you want to find the length of the line. for the first line, the equation for L becomes: L1 = sqrt((a+3)^2 + (2a-1)^2) for the second line, the equation for L becomes: L2 = sqrt((a-8)^2 + (2a+2)^2) simplify both these equations to get: L1 = sqrt(2a^2 + 6a + 9 + 4a^2 - 4a + 1) L2 = sqrt(2a^2 - 16a + 64 + 4a^2 + 8a + 4) simplify these equation further to get: L1 = sqrt(6a^2 + 2a + 10) L2 = sqrt(6a^2 - 8a + 68) set these equations equal to each other to get: sqrt(6a^2 + 2a + 10) = sqrt(6a^2 - 8a + 68 square both sides of this equation to get: 6a^2 + 2a + 10 = 6a^2 - 8a + 68 solve for a to get: a = 5.8 2a = 11.6 those are your solutions for a and 2a and those are the x and y coordinates for point P. your 2 lines are: L1 = the line between the points (-3,1) and (5.8,11.6) L2 = the line between the points (8,-2) and (5.8,11.6) to confirm the lengths of these lines are equal, then solve for the lengths of both lines. you will find that: length of L1 is equal to sqrt(189.8) length of L2 is equal to sqrt(189.8)

OpenStudy (anonymous):

@Athenalexis

OpenStudy (anonymous):

thank you! @Hannah_Waller